OpenStudy (anonymous):

John Gray bought a basic car for $8,250.00, with options that cost $324.00. There's a 6% sales tax in his state and a combined $50.00 license and registration fee. What was John's total cost? A. $14,247.00 B. $9,138.44 C. $9,141.44 D. $10,000.00 I know how to do the problem up to the sales tax part if you could just explain that to me please?!?

OpenStudy (mathmale):

thanks for being so specific about what you want to learn. To calculate the sales tax on the car, convert that 6% to the decimal fraction 0.06; multiply (the price of the car PLUS the price of the options) by 0.06. that's the sales tax amount. Now add together the (price of the car plus options, the sales tax, and the $50.

OpenStudy (anonymous):

one more question what did you mean by multiply (the price of the car PLUS the price of the options) by 0.06? because that is confusing me a little.

OpenStudy (anonymous):

@chandowd

OpenStudy (mathmale):

I'm sorry, orion, but I did reply, but apparently my efforts to post my reply failed. We have to pay tax on the COMBINED

OpenStudy (mathmale):

COST of the car and options. It's easier to add together the two costs and THEN multiply the result by 0.06 to obtain the sales tax.

OpenStudy (anonymous):

is it B.

OpenStudy (anonymous):

i hate this i can't get the correct answer aaarrrrrggggggg!!!!

OpenStudy (anonymous):

i keep coming up with the answer 8814.44 aaarrrggggg!!!

OpenStudy (mathmale):

Car: $8250 Options: $324 Total for care and options: $8574 Sales tax on this total: ($8574)(0.06) = $514.44 Combined fees: $50 Add these together. Why are you so sure you're wrong? In cases like this, please share an image of your work. That way, I could give you good guidance faster.

OpenStudy (anonymous):

8250+324=8574 8574*0.06=514.44 8250+514.44+50=8814.44

OpenStudy (mathmale):

Looks fine to me. Why the aaarrrrgh?

OpenStudy (anonymous):

because it isn't an answer choice so it can't be right!

OpenStudy (anonymous):

help me @mathmale

OpenStudy (mathmale):

What are your answer choices? Is there an $8814? If so, that may be the correct choice, since $8814.44 rounds down to $8814.

OpenStudy (anonymous):

the answer choices are: A. $14,247.00 B. $9,138.44 C. $9,141.44 D. $10,000.00

OpenStudy (anonymous):

@mathmale ^^^^^^^^^^

OpenStudy (anonymous):

it is my last question and i need to turn in my test now please help me out if you can

OpenStudy (anonymous):

@PeaceBrutha help me outs please?!?

OpenStudy (anonymous):

why can't i get this WHY????? I usually can get math easily what is wrong with this question???

OpenStudy (anonymous):

@fritholly1

OpenStudy (anonymous):

Ok, well answers A and D are wrong.

OpenStudy (anonymous):

Correct answer is B Because.... 8250+324=8574 8574*0.06=514.44 8574+514.44+50=9,138.44

OpenStudy (anonymous):

Only problem you were doing is you have to keep the 8,574 with the final equation. Not 8,250.